OpenStudy (anonymous):

Solve the Proportion 4/x = 5/9 4 over x = 5 over 9. @dc104

OpenStudy (anonymous):

4 = 5/9*x 4/(5/9) = x 36/5 = x

OpenStudy (misssunshinexxoxo):

x=36/5

OpenStudy (anonymous):

You have to use cross multiplication 4(9) = 5x Are you with me so far? 36 = 5x 36/5 = x

OpenStudy (studygurl14):

\[\frac{ 4 }{ x } = \frac{ 5 }{ 9 }\] cross-multiply \[4(9) = x(5) \rightarrow 36 = 5x \rightarrow x = 7.2\]

OpenStudy (anonymous):

@StudyGurl14 I am sorry to say but your response is not correct.

OpenStudy (studygurl14):

how come @Geniu

OpenStudy (misssunshinexxoxo):

It is!

OpenStudy (anonymous):

Her response is correct. She just simplified 36/5 try plugging it into a calculator.

OpenStudy (anonymous):

I am not looking for a simplified. I am sorry @StudyGurl14 . I am the one who was wrong here. My textbook said 36/5, it didn't show the simplified version.

OpenStudy (studygurl14):

It's fine @Geniu